# Brain-inspired and Self-based Artificial Intelligence

### v1

@@
Yi Zeng  [^1] These authors contributed equally. [^2] Lead contact Feifei Zhao  [^3] These authors contributed equally. Yuxuan Zhao  [^4] These authors contributed equally. Dongcheng Zhao  [^5] These authors contributed equally. Enmeng Lu [^6] These authors contributed equally. Qian Zhang Yuwei Wang Hui Feng Zhuoya Zhao Jihang Wang Qingqun Kong Yinqian Sun Yang Li Guobin Shen Bing Han Yiting Dong Wenxuan Pan Xiang He Aorigele Bao Jin Wang Brain-inspired Cognitive Intelligence Lab, Institute of Automation, Chinese Academy of Sciences, China. Key Laboratory of Brain Cognition and Brain-inspired Intelligence Technology, Chinese Academy of Sciences, China. School of Artificial Intelligence, University of Chinese Academy of Sciences, China. School of Future Technology, University of Chinese Academy of Sciences, Beijing, China. Department of Philosophy, School of Humanities, University of Chinese Academy of Sciences, Beijing, China
@@ 1 Introduction
In his 1950 paper ”Computing Machinery and Intelligence,” Alan Turing asked the question ”Can machines think?” and introduced the Turing Test to verify whether machines could achieve human-level intelligence Turing (2009). However, Artificial Intelligence (AI) that meets the external requirements of behaviorism is not necessarily ”intelligent” in the true sense, as evidenced by John R. Searle’s Chinese Room thought experiment Searle (1986). As René Descartes said, ”I think, therefore I am,” implying that being aware of one’s thoughts is a crucial element of self-consciousness Descartes and Gilson (1987). However, current AI is merely an appearance of intelligence through information processing, rather than being subjectively aware of oneself, perceive and understand the world from the self perspective and learn from others or perspectives of others like human intelligence. Without the construction and sense of self for AI, we argue there is fundamentally impossible to develop a machine that think in the real sense. Therefore, the concept of Self should play a crucial role in the development of AI.
@@
Some research has already been exploring how to bestow AI and robots with self-awareness akin to humans. Tony J. Prescott argues that, considering the complexity of the Self knowledge, a unified cognitive architecture is indispensable for realizing a multi-faceted self-model Prescott (2020). In fact, given that self-awareness is a vital component of human mind, AI researchers have, for decades, closely aligned their efforts with simulating human minds through the construction of cognitive architectures. Cognitive architectures simulate and understand the human mind through computational modeling, e.g., memory, learning, reasoning, and other cognitive mechanisms and processes. Allen Newell champions that human cognition and mind function as a unified cognitive architecture Newell (1994). Inspired by this, John R. Anderson proposed the Adaptive Control of Thought-Rational (ACT-R) architecture Anderson (2013), a cognitive psychology-based model that integrates the human associative memory model with generative systems architecture, covering perception, memory, decision-making, language, and motor cognitive functions. Allen Newell poured his proposed unified theories of cognition into the Soar cognitive architecture Laird et al. (1987), which was then continuously developed by John E. Laird Laird (2019). Soar is committed to providing the underlying infrastructure that enables the system to perform a variety of cognitive tasks, capable of implementation of a general intelligence. Soar consists of several interacting task-independent modules: short-term and long-term memories, processing modules, learning mechanisms, and interfaces between them.
@@
Unlike the top-down cognitive architecture design of ACT-R, Local, Error-driven and Associative, Biologically Realistic Algorithm (Leabra) O’Reilly (1996) begins from the basic neural mechanisms and integrates them into a unified framework from the bottom to up. Specifically, Leabra incorporates point neuronal computations, and different learning mechanisms organized to correspond to biologically plausible formulations. Then, neurons connect and interact in networks to form small-scale and large-scale brain regional structures. Various brain regional architectures organized and interacted with each other to perform different cognitive functions. SAL Jilk et al. (2008) synthesizes two cognitive architectures, ACT-R and Leabra, treating them as modular and replaceable components, proposing a pluralistic approach reorganized to meet the needs of cognition. Connectionist Learning with Adaptive Rule Induction On-line (CLARION)  Sun (2003) considers the distinction between implicit and explicit cognitive processes and captures their interactions integratively. In addition, Non-Axiomatic Reasoning System (NARS) Wang (1995) shows an experience-based reasoning system using a formal term-oriented language. Semantic Pointer Architecture Unified Network (SPAUN) Eliasmith et al. (2012) has a large-scale spiking neurons organized into subsystems that resemble specific brain regions, capable of performing multiple tasks.
@@
In general, the above cognitive architectures synergize different sub-modules and sub-systems into a unified architecture for accomplishing the full range of cognitive tasks. They approach and simulate human cognitive processes as closely as possible at different levels, covering brain regions such as the prefrontal cortex, basal ganglia, posterior cortex, and hippocampus, performing cognitive tasks such as episodic learning, semantic learning, reinforcement learning, working memory, long-term memory, perception, and action selection. There are also cognitive architectures related to consciousness, such as the Global Workspace Theory Baars (1993), which postulates that attention can transform unconscious processes or modules into conscious awareness. Building on this foundation, Learning Intelligent Distribution Agent (LIDA) Baars and Franklin (2009) focuses on the computational modeling of consciousness and higher-order cognitive control.
@@
To implement BriSe AI, we use an Spiking Neural Network (SNN) Maass (1997) based brain-inspired cognitive intelligence engine (BrainCog) Zeng et al. (2023), seamlessly integrating BrainCog’s components and cognitive functions to build advanced AI models and applications. The basic components of BrainCog include spiking neurons with different levels of granularity, biological learning rules, multiple encoding strategies, and various functional brain region models. Based on these fundamental components, BrainCog implements various learning strategies of spiking neural networks, such as unsupervised learning, supervised learning, associative learning, and reinforcement learning. This supports cognitive functions such as perception and learning, decision-making, motor control, knowledge representation and reasoning, social cognition, etc.
@@ 2 Hierarchical Framework of the Self for AI
The study of self-concept and theories is primarily concentrated in the fields of philosophy and cognitive psychology, while research in the field of artificial intelligence is relatively limited and has not yet achieved complete unity. In 1988, Ulric Neisser proposed five kinds of self-knowledge Neisser (1988), including ecological self, interpersonal self, temporally extended self, private self, and conceptual self. These several ’self’ possess the following characteristics: 1) Ecological self has a point of view and a sense of body ownership. 2) Interpersonal self perceives others as agents similar to oneself and having empathy for them. 3) Temporally extended self is aware of your personal past and future. 4) Private self experiences a stream of consciousness and is aware of your inner life. 5) Conceptual self has a life story, personal goals, motivations, and values. These five kinds of self-knowledge proposed by Ulric Neisser are categorized and generalized from the perspective of the self that humans possess. Building upon this, Tony J. Prescott proposed Robot Self Prescott (2020), which includes the situated self, agential self, spatiotemporal self, interpersonal self, conceptual self, and private self. The main difference from Ulric Neisser’s five kinds of self is the addition of the agential self, emphasizing the characteristics of a robot actively seeking information, selecting actions that generate integrated behavior, and knowing what events it caused in the world.
@@ 3.1 Perception and Learning
BriSe AI’s integrative approach has significantly enhanced its ability to adeptly navigate and address complex challenges across various learning domains, including unsupervised learning Dong et al. (2023), supervised learning Zhao et al. (2023a), few-shot learning Li et al. (2022), continual learning Han et al. (2023), transfer learning He et al. (2023), multi-modal concept learning, and evolutionary learning Shen et al. (2023). These capabilities are crucial for tasks such as advanced visual classification and recognition, where unsupervised and supervised learning play key roles, and few-shot and continual learning strategies are vital in dynamic environments. Particularly, the principles of reinforcement learning have been effectively utilized in complex tasks like robot and drone control, demonstrating adaptability and decision-making prowess in real-world scenarios. Moreover, these achievements in learning and application have also supported the development of higher-level Self within BriSe AI, contributing to its evolving understanding of the Bodily Self, Autonomous Self, Social Self, and Conceptual Self.
@@ 3.2 Bodily Self
Bodily Perception. Bodily perception is one of the core elements of the bodily self. By integrating perceptual information from Level 0 Perception and Learning, intelligent agents engage in multimodal information integration, achieve motor-visual associative learning, construct a body model, and provide the foundation for high-level Self. A central question in bodily perception is how body illusions, such as the rubber hand illusion Botvinick and Cohen (1998) occur. The rubber hand illusion refers to the phenomenon where subjects, when receiving tactile and visual stimuli simultaneously on a rubber hand and their unseen real hand, experience an illusion, perceiving the rubber hand as if it were their own.
@@
We draw insights from biological studies on monkeys and humans and propose a brain-inspired bodily self-perception model Zhao et al. (2023b). This model understands and recognizes one’s own body through the integration and association of multiple somatosensory perceptions and nonproprioception. It enables the robot to pass the mirror test Zeng et al. (2018a) and induce the rubber hand illusion Zhao et al. (2023b). Mirror test is a classic experiment to determine whether animal species have self-awareness. The result of multi-robot mirror self-recognition experiment is shown in Figure  4a. Three robots with identical appearance move randomly in front of the mirror at the same time. The robot can identify which mirror belongs to it by comparing the trajectory predicted according to its own movement with the trajectory actually detected by vision. For example, when the robot is finished moving, the laser light is hit on the left side of the head of the middle robot, and the robot will touch the corresponding position. The robot obtained experimental results similar to those of monkeys in the rubber hand illusion experiment (as shown in Figure  4b). When the visual deflection angle is small, the robot’s behavior decision mainly relies on visual information, and the proprioceptive drift increases with the increase of the visual deflection angle. When the visual deflection angle is medium, the behavior decision of the robot mainly relies on proprioceptive information. As the visual deflection angle increases, the proprioceptive drift increases slowly and tends to be flat. When the visual deflection angle is large, the behavior decision of the robot completely relies on proprioceptive information, and the proprioceptive drift is zero. The robot does not consider the hand in the field of view as its own.
@@
The medial prefrontal cortex (MPFC) and its subregions play a crucial role in self-referential processing D’Argembeau et al. (2007). We explored the anesthesia-induced loss of consciousness on the perception of bodily self based on a primate corticothalamic network model encompassing 29 widely distributed cortical areas cortex (including the medial prefrontal) and the thalamus Zhang et al. (2023). In recent years, anesthetics have emerged as novel instruments for elucidating the neural foundations of consciousness, enabling the precise and controlled simulation of both the loss and regain of consciousness. Propofol, in particular, has gained widespread use as a general anesthetic Alkire et al. (2008). We introduces a brain simulation model including multiple brain regions to explore the neural activity during the propofol-induced loss of consciousness. Our model presents distinct advantages over conventional clinical observation data in that it offers quantifiable and controllable concentrations of propofol, along with high temporal resolution. During general anesthesia, bodily perception and consciousness diminish over time, correlating with the progressive involvement of different brain areas in rhythm propagation. This spatial distribution is distinct and systematic (as shown in Figure 5), with a noticeable transition of rhythmic activity from the occipital lobe to the prefrontal and then to the frontal lobe across the first three-time groups. Ultimately, a complete cortical diffusion occurs. This model-generated process aligns with the clinically observed phenomenon of $\alpha$-rhythm anteriorization Vijayan et al. (2013).
@@ 3.3 Autonomous Self
Classical Conditioning. Classical conditioning is one of the most fundamental and crucial learning mechanisms in the Autonomous Self. It aids organisms in adapting to the complex and dynamic natural environment, allowing them to anticipate and adapt to environmental changes predictably, thereby enhancing their survival capabilities. One of the most renowned experiments in classical conditioning is conducted by Pavlov Pavlov (1927). In this experiment, when dogs are presented with food (unconditioned stimulus, $US$), they exhibit salivation (unconditioned response, $UR$). In Pavlov’s research, a sound (conditioned stimulus, $CS$) is introduced before each instance of food presentation. After numerous trials, the dogs start salivating upon hearing the tone (conditioned response, $CR$).
Based on existing research findings in disciplines such as biology and neuroscience in the field of classical conditioning, we proposed a brain-inspired classical conditioning model Zhao et al. (2021). This model integrates the biological research findings that have reached consensus in the field of classical conditioning into a spiking neural network. This model revolves around the interpositus nucleus (IPN), pontine nuclei (PN), granule cell (GC), Purkinje cell (PU), parallel fiber, climbing fiber, and inferior olive. It is designed to create pathways for conditioned reflexes and unconditioned reflexes. The synaptic plasticity mechanisms based on associative learning occur between PN and IPN, as well as between GC and PU. Compared to other computational models, the brain-inspired classical conditioning model can reproduce up to 15 classic conditioned phenomena and provide reasonable explanations from a computational perspective, which helps to reveal the biological mechanism by which organisms establish classical conditioning. In addition, the model can be deployed on robots, enabling them to establish classical conditioning similar to organisms. After verification, it has been demonstrated that this model endows robots with the capability of speed generalization. The speed generalization ability refers to the ability of robots to complete navigation tasks through conditional reflection at lower speeds and to complete navigation tasks without training at higher speeds. In a simulation environment with precise and controllable speed, the robot can adapt up to 3.5 times the speed and complete navigation tasks.
@@
To realize decision-making function, we designed the multi-brain region coordinated spiking neural network to model the decision-making neural circuit of the mammalian brain (prefrontal cortex-basal ganglia-thalamus-premotor cortex circuit) Zhao et al. (2018a). The excitatory and inhibitory connections between multiple brain regions are optimized by biologically plausible synaptic plasticity, where dopamine-modulated synaptic plasticity plays a key role in achieving trial-and-error learning. The decision-making process comprises subsystems including the agent’s perception of the environment, winner-takes-all action selection, environmental feedback, strategy adjustment, and autonomous exploration. It collaboratively incorporates reinforcement learning strategies to acquire skills and accomplish tasks. As shown in Figure 6, the proposed model supports agents playing games Sun et al. (2022), UAV flying through doors and windows Zhao et al. (2018b), obstacle avoidance in real-world scenarios Zhao et al. (2018a), and the safe flight of multiple drones within a limited area Zhao et al. (2022).

### v2

@@
Yi Zeng  [^1] These authors contributed equally. [^2] Lead contact Feifei Zhao  [^3] These authors contributed equally. Yuxuan Zhao  [^4] These authors contributed equally. Dongcheng Zhao  [^5] These authors contributed equally. Enmeng Lu [^6] These authors contributed equally. Qian Zhang Yuwei Wang Hui Feng Zhuoya Zhao Jihang Wang Qingqun Kong Yinqian Sun Yang Li Guobin Shen Bing Han Yiting Dong Wenxuan Pan Xiang He Aorigele Bao Jin Wang Brain-inspired Cognitive AI Lab, Institute of Automation, Chinese Academy of Sciences, China. Beijing Key Laboratory of Safe AI and Superalignment, China. Beijing Institute of AI Safety and Governance, China. Key Laboratory of Brain Cognition and Brain-inspired Intelligence Technology, Chinese Academy of Sciences, China. University of Chinese Academy of Sciences, China. Department of Philosophy, School of Humanities, University of Chinese Academy of Sciences, China. Long-term AI, China.
@@ 1 Introduction
In his 1950 paper ”Computing Machinery and Intelligence,” Alan Turing asked the question ”Can machines think?” and introduced the Turing Test to verify whether machines could achieve human-level intelligence [1]. However, Artificial Intelligence (AI) that meets the external requirements of behaviorism is not necessarily ”intelligent” in the true sense, as evidenced by John R. Searle’s Chinese Room thought experiment [2]. As René Descartes said, ”I think, therefore I am,” implying that being aware of one’s thoughts is a crucial element of self-consciousness [3]. However, current AI is merely an appearance of intelligence through information processing, rather than being subjectively aware of oneself, perceive and understand the world from the self perspective and learn from others or perspectives of others like human intelligence. Without the construction and sense of self for AI, we argue there is fundamentally impossible to develop a machine that think in the real sense. Therefore, the concept of Self should play a crucial role in the development of AI.
@@
Some research has already been exploring how to bestow AI and robots with self-awareness akin to humans. Tony J. Prescott argues that, considering the complexity of the Self knowledge, a unified cognitive architecture is indispensable for realizing a multi-faceted self-model [4]. In fact, given that self-awareness is a vital component of human mind, AI researchers have, for decades, closely aligned their efforts with simulating human minds through the construction of cognitive architectures. Cognitive architectures simulate and understand the human mind through computational modeling, e.g., memory, learning, reasoning, and other cognitive mechanisms and processes. Allen Newell champions that human cognition and mind function as a unified cognitive architecture [5]. Inspired by this, John R. Anderson proposed the Adaptive Control of Thought-Rational (ACT-R) architecture [6], a cognitive psychology-based model that integrates the human associative memory model with generative systems architecture, covering perception, memory, decision-making, language, and motor cognitive functions. Allen Newell poured his proposed unified theories of cognition into the Soar cognitive architecture [7], which was then continuously developed by John E. Laird [8]. Soar is committed to providing the underlying infrastructure that enables the system to perform a variety of cognitive tasks, capable of implementation of a general intelligence. Soar consists of several interacting task-independent modules: short-term and long-term memories, processing modules, learning mechanisms, and interfaces between them.
@@
Unlike the top-down cognitive architecture design of ACT-R, Local, Error-driven and Associative, Biologically Realistic Algorithm (Leabra) [9] begins from the basic neural mechanisms and integrates them into a unified framework from the bottom to up. Specifically, Leabra incorporates point neuronal computations, and different learning mechanisms organized to correspond to biologically plausible formulations. Then, neurons connect and interact in networks to form small-scale and large-scale brain regional structures. Various brain regional architectures organized and interacted with each other to perform different cognitive functions. SAL [10] synthesizes two cognitive architectures, ACT-R and Leabra, treating them as modular and replaceable components, proposing a pluralistic approach reorganized to meet the needs of cognition. Connectionist Learning with Adaptive Rule Induction On-line (CLARION)  [11] considers the distinction between implicit and explicit cognitive processes and captures their interactions integratively. In addition, Non-Axiomatic Reasoning System (NARS) [12] shows an experience-based reasoning system using a formal term-oriented language. Semantic Pointer Architecture Unified Network (SPAUN) [13] has a large-scale spiking neurons organized into subsystems that resemble specific brain regions, capable of performing multiple tasks.
@@
In general, the above cognitive architectures synergize different sub-modules and sub-systems into a unified architecture for accomplishing the full range of cognitive tasks. They approach and simulate human cognitive processes as closely as possible at different levels, covering brain regions such as the prefrontal cortex, basal ganglia, posterior cortex, and hippocampus, performing cognitive tasks such as episodic learning, semantic learning, reinforcement learning, working memory, long-term memory, perception, and action selection. There are also cognitive architectures related to consciousness, such as the Global Workspace Theory [14], which postulates that attention can transform unconscious processes or modules into conscious awareness. Building on this foundation, Learning Intelligent Distribution Agent (LIDA) [15] focuses on the computational modeling of consciousness and higher-order cognitive control.
@@
To implement BriSe AI, we use an Spiking Neural Network (SNN) [16] based brain-inspired cognitive intelligence engine (BrainCog) [17], seamlessly integrating BrainCog’s components and cognitive functions to build advanced AI models and applications. The basic components of BrainCog include spiking neurons with different levels of granularity, biological learning rules, multiple encoding strategies, and various functional brain region models. Based on these fundamental components, BrainCog implements various learning strategies of spiking neural networks, such as unsupervised learning, supervised learning, associative learning, and reinforcement learning. This supports cognitive functions such as perception and learning, decision-making, motor control, knowledge representation and reasoning, social cognition, etc.
@@ 2 Hierarchical Framework of the Self for AI
The study of self-concept and theories is primarily concentrated in the fields of philosophy and cognitive psychology, while research in the field of artificial intelligence is relatively limited and has not yet achieved complete unity. In 1988, Ulric Neisser proposed five kinds of self-knowledge [18], including ecological self, interpersonal self, temporally extended self, private self, and conceptual self. These several ’self’ possess the following characteristics: 1) Ecological self has a point of view and a sense of body ownership. 2) Interpersonal self perceives others as agents similar to oneself and having empathy for them. 3) Temporally extended self is aware of your personal past and future. 4) Private self experiences a stream of consciousness and is aware of your inner life. 5) Conceptual self has a life story, personal goals, motivations, and values. These five kinds of self-knowledge proposed by Ulric Neisser are categorized and generalized from the perspective of the self that humans possess. Building upon this, Tony J. Prescott proposed Robot Self [4], which includes the situated self, agential self, spatiotemporal self, interpersonal self, conceptual self, and private self. The main difference from Ulric Neisser’s five kinds of self is the addition of the agential self, emphasizing the characteristics of a robot actively seeking information, selecting actions that generate integrated behavior, and knowing what events it caused in the world.
@@ 3.1 Perception and Learning
BriSe AI’s integrative approach has significantly enhanced its ability to adeptly navigate and address complex challenges across various learning domains, including unsupervised learning [19], supervised learning [20], few-shot learning [21], continual learning [22], transfer learning [23], multi-modal concept learning, and evolutionary learning [24]. These capabilities are crucial for tasks such as advanced visual classification and recognition, where unsupervised and supervised learning play key roles, and few-shot and continual learning strategies are vital in dynamic environments. Particularly, the principles of reinforcement learning have been effectively utilized in complex tasks like robot and drone control, demonstrating adaptability and decision-making prowess in real-world scenarios. Moreover, these achievements in learning and application have also supported the development of higher-level Self within BriSe AI, contributing to its evolving understanding of the Bodily Self, Autonomous Self, Social Self, and Conceptual Self.
@@ 3.2 Bodily Self
Bodily Perception. Bodily perception is one of the core elements of the bodily self. By integrating perceptual information from Level 0 Perception and Learning, intelligent agents engage in multimodal information integration, achieve motor-visual associative learning, construct a body model, and provide the foundation for high-level Self. A central question in bodily perception is how body illusions, such as the rubber hand illusion [25] occur. The rubber hand illusion refers to the phenomenon where subjects, when receiving tactile and visual stimuli simultaneously on a rubber hand and their unseen real hand, experience an illusion, perceiving the rubber hand as if it were their own.
@@
We draw insights from biological studies on monkeys and humans and propose a brain-inspired bodily self-perception model [26]. This model understands and recognizes one’s own body through the integration and association of multiple somatosensory perceptions and nonproprioception. It enables the robot to pass the mirror test [27] and induce the rubber hand illusion [26]. Mirror test is a classic experiment to determine whether animal species have self-awareness. The result of multi-robot mirror self-recognition experiment is shown in Figure  4a. Three robots with identical appearance move randomly in front of the mirror at the same time. The robot can identify which mirror belongs to it by comparing the trajectory predicted according to its own movement with the trajectory actually detected by vision. For example, when the robot is finished moving, the laser light is hit on the left side of the head of the middle robot, and the robot will touch the corresponding position. The robot obtained experimental results similar to those of monkeys in the rubber hand illusion experiment (as shown in Figure  4b). When the visual deflection angle is small, the robot’s behavior decision mainly relies on visual information, and the proprioceptive drift increases with the increase of the visual deflection angle. When the visual deflection angle is medium, the behavior decision of the robot mainly relies on proprioceptive information. As the visual deflection angle increases, the proprioceptive drift increases slowly and tends to be flat. When the visual deflection angle is large, the behavior decision of the robot completely relies on proprioceptive information, and the proprioceptive drift is zero. The robot does not consider the hand in the field of view as its own.
@@
The medial prefrontal cortex (MPFC) and its subregions play a crucial role in self-referential processing [28]. We explored the anesthesia-induced loss of consciousness on the perception of bodily self based on a primate corticothalamic network model encompassing 29 widely distributed cortical areas cortex (including the medial prefrontal) and the thalamus [29]. In recent years, anesthetics have emerged as novel instruments for elucidating the neural foundations of consciousness, enabling the precise and controlled simulation of both the loss and regain of consciousness. Propofol, in particular, has gained widespread use as a general anesthetic [30]. We introduces a brain simulation model including multiple brain regions to explore the neural activity during the propofol-induced loss of consciousness. Our model presents distinct advantages over conventional clinical observation data in that it offers quantifiable and controllable concentrations of propofol, along with high temporal resolution. During general anesthesia, bodily perception and consciousness diminish over time, correlating with the progressive involvement of different brain areas in rhythm propagation. This spatial distribution is distinct and systematic (as shown in Figure 5), with a noticeable transition of rhythmic activity from the occipital lobe to the prefrontal and then to the frontal lobe across the first three-time groups. Ultimately, a complete cortical diffusion occurs. This model-generated process aligns with the clinically observed phenomenon of $\alpha$-rhythm anteriorization [31].
@@ 3.3 Autonomous Self
Classical Conditioning. Classical conditioning is one of the most fundamental and crucial learning mechanisms in the Autonomous Self. It aids organisms in adapting to the complex and dynamic natural environment, allowing them to anticipate and adapt to environmental changes predictably, thereby enhancing their survival capabilities. One of the most renowned experiments in classical conditioning is conducted by Pavlov [32]. In this experiment, when dogs are presented with food (unconditioned stimulus, $US$), they exhibit salivation (unconditioned response, $UR$). In Pavlov’s research, a sound (conditioned stimulus, $CS$) is introduced before each instance of food presentation. After numerous trials, the dogs start salivating upon hearing the tone (conditioned response, $CR$).
Based on existing research findings in disciplines such as biology and neuroscience in the field of classical conditioning, we proposed a brain-inspired classical conditioning model [33]. This model integrates the biological research findings that have reached consensus in the field of classical conditioning into a spiking neural network. This model revolves around the interpositus nucleus (IPN), pontine nuclei (PN), granule cell (GC), Purkinje cell (PU), parallel fiber, climbing fiber, and inferior olive. It is designed to create pathways for conditioned reflexes and unconditioned reflexes. The synaptic plasticity mechanisms based on associative learning occur between PN and IPN, as well as between GC and PU. Compared to other computational models, the brain-inspired classical conditioning model can reproduce up to 15 classic conditioned phenomena and provide reasonable explanations from a computational perspective, which helps to reveal the biological mechanism by which organisms establish classical conditioning. In addition, the model can be deployed on robots, enabling them to establish classical conditioning similar to organisms. After verification, it has been demonstrated that this model endows robots with the capability of speed generalization. The speed generalization ability refers to the ability of robots to complete navigation tasks through conditional reflection at lower speeds and to complete navigation tasks without training at higher speeds. In a simulation environment with precise and controllable speed, the robot can adapt up to 3.5 times the speed and complete navigation tasks.
@@
To realize decision-making function, we designed the multi-brain region coordinated spiking neural network to model the decision-making neural circuit of the mammalian brain (prefrontal cortex-basal ganglia-thalamus-premotor cortex circuit) [34]. The excitatory and inhibitory connections between multiple brain regions are optimized by biologically plausible synaptic plasticity, where dopamine-modulated synaptic plasticity plays a key role in achieving trial-and-error learning. The decision-making process comprises subsystems including the agent’s perception of the environment, winner-takes-all action selection, environmental feedback, strategy adjustment, and autonomous exploration. It collaboratively incorporates reinforcement learning strategies to acquire skills and accomplish tasks. As shown in Figure 6, the proposed model supports agents playing games [35], UAV flying through doors and windows [36], obstacle avoidance in real-world scenarios [34], and the safe flight of multiple drones within a limited area [37].